An electron and a proton separated by a distance of 1.5 × 10 ^ −10 m. Calculate the

magnitude of the electrostatic force between the particles.


1
Expert's answer
2021-07-06T08:23:26-0400

We can find the electrostatic force between the electron and proton as follows:


"F=\dfrac{kq_eq_p}{r^2},""F=\dfrac{9\cdot10^9\ \dfrac{N\cdot m^2}{C^2}\cdot1.6\cdot10^{-19}\ C\cdot1.6\cdot10^{-19}\ C}{(1.5\cdot10^{-10}\ m)^2},""F=1.024\cdot10^{-8}\ N."